OpenStudy (anonymous):

Integrate:

OpenStudy (anonymous):

\[\int\limits{\sqrt{4x+1}} dx \]

OpenStudy (anonymous):

the answer is \[(1/6)[(4x+1)^{3/2}]\] set u = 4x + 1 and du = 4dx \\[1/4\int\limits\limits\limits_{}^{}(u ^{1/2)^{}}du\]

OpenStudy (anonymous):

cheers